Alex bounced a 0.59 kg rubber ball off the floor. The ball's speed just before impacting the floor was 12 m/s, and the speed just after the impact was 4.5 m/s. What implies Ly was applied to the ball by the floor? If the ball was in contact with the flow for 0.036s, what was the magnitude of the average force applied by the floor on the ball?

Find Impluse

Find Average force
Answer #1

m =0.59 kg , u =12 m/s, v =4.5 m/s, t =0.036 s

Impulse = change in momentum = mv- mu = 0.59(4.5-12)

= -4.425 kg.m/s

Average force = impulse/time = 4.425/0.036 = -123 N